OpenStudy (anonymous):

Which equation is not equivalent to the formula e = mc? m equals e over c c equals e over m e = cm m equals c over e

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

e=mc divide both sides by c \[\frac{ e }{ c }=\frac{ mc }{ c }=>\frac{ e }{ c }=m\]so clearly we can see that 4th one is wrong.

OpenStudy (michele_laino):

Hint: let's start from m=c over e, then i multiply botth sides by e, so I can write this: \[m \cdot e = \frac{c}{e} \cdot e\]

OpenStudy (anonymous):

wait .. so d is the right answer?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

ok ty both

OpenStudy (michele_laino):

:)